# On the contraction properties of Sinkhorn semigroups

### v1

@@ 5.1 Strong mixing models
Proof.
$(\lambda_{U},\pi_{2n+1})$ $=$ $(\pi_{2n}{\cal S}_{2n+1},\nu_{V}{\cal S}_{2n+1}),$
Corollary 5.2.
Several examples satisfying condition $\jmath({\cal Q})>0$ are discussed in Section 2.4 in [35]. For instance, consider a parameter $\delta>0$ and drift function $x\mapsto b(x)$ with bounded oscillations from $\mathbb{X}=\mathbb{R}^{d}$ into itself. The choice ${\cal Q}(x,dy)~{}\propto~{}e^{-\delta\,\|y-b(x)\|}~{}dx$ yields
In continuous time settings, condition $\jmath({\cal Q})>0$ is also satisfied for Markov transitions of elliptic diffusions on compact manifolds $\mathbb{X}$ –see for instance the pioneering work of Aronson [7], Nash [64] and Varopoulos [78] on Gaussian estimates for heat kernels on manifolds. Nevertheless, the above condition is not met for linear Gaussian transitions.
@@ 5.2 Linear-Gaussian models
[…]
@@
$(m,\overline{m})\in(\mathbb{R}^{d}\times\mathbb{R}^{d})\quad\text{and}\quad(\sigma,\overline{\sigma})\in({\cal S}^{+}_{d}\times{\cal S}^{+}_{d}).$
@@
for some given $(\alpha,\beta,\tau)\in\left(\mathbb{R}^{d}\times{\cal G}l_{d}\times{\cal S}^{+}_{d}\right)$. In this context, we have $\jmath({\cal Q})=0$ and Hilbert projective techniques cannot be applied. However, a refined exponential stability analysis of the Sinkhorn algorithm for Gaussian models has been developed in [2]. For instance, following Remark 5.4 in [2], there exists some $n_{0}\geq 1$ such that for any $n\geq n_{0}$ we have the estimate
@@
A proof is provided in Appendix A (on page A).
@@
$\varphi(x)$ ::=:= $\exp{\left(\delta\,((x-m)^{\prime}\sigma^{-1}(x-m))^{1/2}\right)}\quad\text{and}$
$\psi(y)$ ::=:= $\exp{\left(\delta\,((y-\overline{m})^{\prime}\,\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}\right)}.$ (55)
@@
$\varphi(x)$ ::=:= $\exp{\left(\delta\,(x-m)^{\prime}\sigma^{-1}(x-m)\right)}\quad\text{and}$
$\psi(y)$ ::=:= $\exp{\left(\delta\,(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m})\right)}.$ (56)
A detailed proof of the above assertions is provided in Appendix A on page A.
@@ 5.3 Finite mixture models
$-U_{i}(x)=\log{g_{\sigma_{i}}(x-m_{i})}\quad\mbox{\rm and}\quad-V_{j}(y):=\log{g_{\overline{\sigma}_{j}}(y-\overline{m}_{j})},$ (59)
@@
$\lambda_{U}(dx):=\frac{1}{N}\sum_{1\leq i\leq N}g_{\sigma}(x-X_{i})~{}dx.$ (61)
is a well known Gaussian mixture approximation of the distribution $\eta(dx)$ (as the number of samples $N\rightarrow\infty$). The covariance matrix $\sigma$ is a bandwidth parameter that controls the amount and orientation of the smoothing. Note that (61) coincides with (58) when $(m_{i},\sigma_{i})=(X_{i},\sigma)$ with the counting measure $\xi_{0}(i)=1/N$ on the index set $E_{0}=\{1,\ldots,N\}$. For a more detailed discussion on these non parametric statistical estimation techniques we refer to the pioneering articles [67, 69] as well as the books [16, 73, 79].
The Gaussian mixture distribution (61) can be used to estimate complex and unknown data distributions arising in generative models, see for instance [31] and references therein. Choosing $(\alpha,\beta)=(0,I)$ and $\tau=tI$; for some $t>0$ in (51), the Markov transition ${\cal Q}(x,dy)$ coincides with the the transition semigroup of the $d$-dimensional Brownian motion, namely
@@
[…]
@@ Appendix A The Gaussian Sinkhorn algorithm
and $\upsilon_{0}:=\overline{\sigma}^{-1/2}\tau~{}\overline{\sigma}^{-1/2}$, then it can be shown (see Theorem 4.3 in [2]) that the recursions
@@
${\cal Z}^{\circ}_{2n}(y)$ ::=:= $\overline{m}+\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~{}G$
${\cal Z}^{\circ}_{2n+1}(x)$ ::=:= $m+\beta_{2n+1}^{\circ}(x-m)+\left(\tau_{2n+1}^{\circ}\right)^{1/2}~{}G$
@@
Moreover, Corollary 4.4 in [2] yields the uniform estimates
@@
$\begin{array}[]{l}\displaystyle\frac{1}{\psi(y)}~{}\mathbb{E}\left(\psi\left({\cal Z}^{\circ}_{2n}(y)\right)\right)\\ \\ \leq\displaystyle\frac{1}{\sqrt{\mbox{\rm det}(I-\delta~{}I)}}~{}\exp{\left(-\frac{\delta}{2}~{}\left(1-\frac{1}{1-\delta}(1+\lambda_{\sf min}(\varpi))^{-2}\right)~{}(y-\overline{m})^{\prime}~{}\overline{\sigma}^{-1}~{}(y-\overline{m})\right)}.\end{array}$
@@ Appendix B Hilbert metric contractions
We recall the Hilbert contraction coefficient formula ([9, 11, 12, 50], see also [8, 22, 59])
[…]
@@
[…]
@@ Appendix C Weighted norm contractions
$\begin{array}[]{ccccc}\varepsilon&:&[r_{0},\infty[&\mapsto&]0,1]\\ &&r&\leadsto&\varepsilon(r)\\ \end{array}$
@@
When $\varphi\in{\cal B}_{\infty}(\mathbb{X})$ the $r$-sub-level sets $C_{\varphi}(r)$ are compact condition (77) is met as soon as
@@
[…]
Theorem C.1 is a consequence of Lemma 2.3 in [36], see also Theorem 8.2.21 in [38] and Lemma 2.1 in [6]. Several examples satisfying the Lyapunov condition (75) and the local minorization conditions (82) are discussed in [6, 38].
@@
[…]
@@ Appendix D Schrödinger potential functions
D.1 Some quantitative estimates
@@
Observe that
$\text{if}\quad\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}<\infty\quad\text{then}\quad e^{\delta U}=\frac{1}{e^{-\delta U}}\geq\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}>0$
and this yields
@@
We set $\sigma_{\beta}=\beta\sigma\beta^{\prime}$ and $\beta_{\tau}:=\tau^{-1}\beta$. With this notation at hand, after some elementary computations (cf. for instance Exemple 6.6 in [2]) we have
@@ D.2 Lyapunov inequalities
${\cal S}_{2n}\left(\psi\right)/\varphi\leq c_{\delta}~{}e^{-{\cal U}_{\delta}}~{}\in{\cal B}_{0}(\mathbb{X})\quad\mbox{and}\quad{\cal S}_{2n+1}\left(\varphi\right)/\psi\leq c_{\delta}~{}e^{-{\cal V}_{\delta}}~{}\in{\cal B}_{0}(\mathbb{Y})$ (96)
@@
D.3 Proof of Theorem 4.3
Following exactly the same argument leading to (81) in Appendix C, the Lyapunov inequalities (96) in Proposition D.9 ensure that the drift conditions stated in (39) hold with with $n_{0}=1$ and $(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})$.
For $\lambda$-almost any $x\in C_{\varphi}(r)$ we have
@@
$\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\int~{}e^{W_{U}(y)}~{}\nu_{V}(dy)\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\nu_{(1-\delta)V}\left(e^{W_{U}}\right)$
@@
$\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\lambda_{U}(e^{W^{V}})\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)$
@@
On the other hand, using Lemma D.8 for any $x\in C_{\varphi}(r)$ we arrive at
@@
Note that for any $r>1$ we have
@@
and this yields, for any $x\in C_{\varphi}(r)$, the estimate
@@
$\left|\frac{d\pi_{2n}}{d\nu_{V}}(y)-1\right|\leq\|\frac{d\pi_{2p}}{d\nu_{V}}\|_{\psi}~{}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|(\delta_{y}-\nu_{V})\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\right)\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\psi}.$
@@
$~{}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|(\delta_{y}-\nu_{V})\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\right)\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\psi}$$\leq$ $c_{\kappa,1}~{}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|(\delta_{y}-\nu_{V})\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\right)\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\psi_{\kappa}}$
@@
where the last assertion comes from the fact that $\log{(1+u)}\leq u$, for any $01+u>$ 0. This implies that

### v2

@@ 5.6 Exponential marginal models
[…]
which yields the cost function
$W(x,y)\leq c_{w}+\varsigma|x-y|\leq c_{w}+\varsigma~x+\varsigma~y,\quad\mbox{\rm with $c_{w}:=\log{(2/\varsigma)}$ }.$
In this context, condition ${\cal H}^{\prime}_{\delta}(U,V)$ holds for any $({\varsigma}/{\tau_{u}})\vee({\varsigma}/{\tau_{v}})<\delta<{1}/{2}$.
5.7 Semi-compact models
In many practical situations of interest, the target state spaces are different. Next we illustrate this situation with a semi-discrete model and a semi-compact models.
Semi-discrete model
Consider a finite or countable subset $\mathbb{X}\subset\mathbb{R}^{d}$ equipped with the counting measure $\lambda$ and the discrete topology. In this context, the potential function $U$ is a function
$U~:~x\in\mathbb{X}\mapsto U(x)\in[0,\infty[\quad\mbox{\rm such that}\quad\sum_{x\in\mathbb{X}}e^{-U(x)}=1.$
Let $\nu_{V}(dy)=e^{-V(y)}~dy$ be a Boltzmann-Gibbs probability measure on $\mathbb{R}^{d}$ and assume that $(U,V,W)$ are continuous functions satisfying the polynomial growth conditions (83). In this context, the sub-level sets of $U$ are finite, thus compact in $\mathbb{X}$.
Following the proof of Proposition 5.1 these conditions ensure that ${\cal H}^{\prime}_{\delta}(U,V)$ holds for some $0<\delta<1/2$ as soon as condition (84) is satisfied. The analysis of heavy tailed transitions follows word-for-word the same lines of arguments as in Section 5.2, this it is skipped.
Semi-compact model
Consider the uniform measure $\lambda_{U}$ on a compact space $\mathbb{X}\subset\mathbb{Y}=\mathbb{R}^{d}$ associated with a null potential $U=0$, which yields $\lambda_{U}=\lambda$. Let $\nu_{V}$ be a centered Gaussian distribution with unit variance on $\mathbb{Y}$ and assume that the transition potential is chosen so that
$W(x,y)\leq w(x)+a~\|y\|^{p}\Longrightarrow W_{U}(y)\leq\lambda(w)+b~\|y\|^{p}$
for some continuous function $w(x)$ and some parameters $a\leq 2$ and $p\leq 1$. These models satisfy ${\cal H}^{\prime}_{\delta}(U,V)$ as soon as
$(p=2~~\mbox{\rm and}~~b<\delta/2)\quad\mbox{\rm or}\quad(p<2)$ (86)
We check this claim using the fact that
[…]
@@ 5.8 Linear-Gaussian models
Consider the state spaces $\mathbb{X}=\mathbb{Y}=\mathbb{R}^{d}$, for some integer $d\geq 1$ and the Lebesgue measure $\lambda(dz)=\nu(dz)=dz$. Hereafter, points $x$ in the Euclidean space $\mathbb{X}=\mathbb{R}^{d}$ are represented by $d$-dimensional column vectors (or, equivalently, by $d\times 1$ matrices). When there is no possible confusion, we use the notation $\|\cdot\|$ for any equivalent matrix or vector norm.
@@
[…]
@@
[…]
@@
A proof is provided in the Appendix A (on page A).
@@
$\varphi(x)$ $:=$ $\exp{\left(\delta\,((x-m)^{\prime}\sigma^{-1}(x-m))^{1/2}\right)}\quad\text{and}$
$\psi(y)$ $:=$ $\exp{\left(\delta\,((y-\overline{m})^{\prime}\,\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}\right)}.$ (92)
@@
$\varphi(x)$ $:=$ $\exp{\left(\delta\,(x-m)^{\prime}\sigma^{-1}(x-m)\right)}\quad\text{and}$
$\psi(y)$ $:=$ $\exp{\left(\delta\,(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m})\right)}.$ (93)
A detailed proof of the above assertions is provided in Appendix A (on page A).
@@ 5.9 Finite mixture models
$U_{i}(x)=\log{g_{\sigma_{i}}(x-m_{i})}\quad\mbox{\rm and}\quad-V_{j}(y):=\log{g_{\overline{\sigma}_{j}}(y-\overline{m}_{j})},$ (96)
@@
$\lambda_{U}(dx):=\frac{1}{N}\sum_{1\leq i\leq N}g_{\sigma}(x-X_{i})~dx$ (98)
is a well known Gaussian mixture approximation of the distribution $\eta(dx)$ (as the number of samples $N\rightarrow\infty$). The covariance matrix $\sigma$ is a bandwidth parameter that controls the amount and orientation of the smoothing. Note that (98) coincides with (95) when $(m_{i},\sigma_{i})=(X_{i},\sigma)$ with the counting measure $\xi_{0}(i)=1/N$ on the index set $E_{0}=\{1,\ldots,N\}$. For a more detailed discussion on these non parametric statistical estimation techniques we refer to the pioneering articles [parzen, rosenblatt] as well as the books [chacon, simonov, wand].
The Gaussian mixture distribution (98) can be used to estimate complex and unknown data distributions arising in generative models, see for instance [doucet-bortoli] and references therein.
Remark 5.6.
Schrödinger bridges and the Sinkhorn algorithm are often used in practice to solve high dimensional optimal transport problems. In our context, choosing $\tau=tI$ and $(\alpha,\beta)=(0,I)$ in (88), the Schrödinger bridge in the limit $t\rightarrow 0$ of vanishing regularization solves the Monge-Kantorovich problem and converge towards the optimal $2$-Wasserstein coupling. In this case, ${\cal S}_{0}=Q$ coincides with the heat kernel on a time horizon $t$ (see also Remark 2.3 in [adm-24]), namely,
@@
In many cases, the Markov transition ${\cal Q}(x,dy)$ of interest is generated by a stochastic diffusion between time 0 and a time horizon $t$ –see, e.g., [leonard, stromme2023sampling, kato2024large]. In our context, if we choose $(\alpha,\beta)=(0,e^{tA})$ and $\tau=\tau_{t}:=\int_{0}^{t}e^{sA}e^{sA^{\prime}}ds$ for some Hurwitz matrix $A$ and some $t>0$, ${\cal Q}(x,dy)$ coincides with the transition semigroup of the Ornstein-Uhlenbeck diffusion (see Remark 2.3 in [adm-24]). In this case, the l.h.s. condition in (97) applied to (98) is met for any $t\geq t_{0}$ provided that
@@ Appendix A The Gaussian Sinkhorn algorithm
and $\upsilon_{0}:=\overline{\sigma}^{-1/2}\tau~\overline{\sigma}^{-1/2}$, then it can be shown (see Theorem 4.3 in [adm-24]) that the recursions
@@
${\cal Z}^{\circ}_{2n}(y)$ $:=$ $\overline{m}+\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~G$
${\cal Z}^{\circ}_{2n+1}(x)$ $:=$ $m+\beta_{2n+1}^{\circ}(x-m)+\left(\tau_{2n+1}^{\circ}\right)^{1/2}~G$
@@
Moreover, Corollary 4.4 in [adm-24] yields the uniform estimates
@@
$\frac{1}{\psi(y)}~\mathbb{E}\left(\psi\left({\cal Z}^{\circ}_{2n}(y)\right)\right)$ $\leq$ $\frac{1}{\sqrt{\mbox{\rm det}(I-\delta~I)}}$$\times$
$\times~\exp{\left(-\frac{\delta}{2}~\left(1-\frac{1}{1-\delta}(1+\lambda_{\sf min}(\varpi))^{-2}\right)~(y-\overline{m})^{\prime}~\overline{\sigma}^{-1}~(y-\overline{m})\right)}.$
@@ Appendix B Hilbert metric contractions
We recall the Hilbert contraction coefficient formula ([bauer, birkhoff, birkhoff-2, hopf], see also [atar, cohen, legland])
$\raisebox{2.0pt}{$\chi$}_{H}({\cal K})=\tanh{\left(\frac{1}{4}\log\sup_{(x_{1},x_{2})\in\mathbb{X}^{2}}H(\delta_{x_{1}}{\cal K},\delta_{x_{2}}{\cal K})\right)}.$ (110)
@@
$\raisebox{2.0pt}{$\chi$}_{H}({\cal K})=\frac{1-\sqrt{\hbar({\cal K})}}{1+\sqrt{\hbar({\cal K})}}.$ (111)
@@ Appendix C Weighted norm contractions
[…]
@@
When $\varphi\in{\cal B}_{\infty}(\mathbb{X})$ the $r$-sub-level sets $C_{\varphi}(r)$ are compact and condition (114) is met as soon as we have
@@
$\raisebox{2.0pt}{$\chi$}_{{\varphi_{\kappa},\psi_{\kappa}}}({\cal K})\leq\rho\quad\mbox{with}\quad(\varphi_{\kappa},\psi_{\kappa}):=(1+\kappa\varphi,1+\kappa\psi).$ (119)
Theorem C.2 is a consequence of Lemma 2.3 in [horton-dp], see also Theorem 8.2.21 in [penev] and Lemma 2.1 in [dpa]. Several examples satisfying the Lyapunov condition (112) and the local minorization conditions (119) are discussed in [dpa, penev].
@@
We check that if Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied for some integer $n_{0}\geq 0$, then (53) and (54) also hold for any $n>n_{0}$.
@@
$\raisebox{2.0pt}{$\chi$}_{{\varphi_{\kappa},\psi_{\kappa}}}({\cal S}_{2n})\vee\raisebox{2.0pt}{$\chi$}_{{\psi_{\kappa},\varphi_{\kappa}}}({\cal S}_{2n+1})\leq\rho.$
@@ Appendix D Schrödinger potential functions
D.1 Commutation formulae
For any $n\geq 0$ we have the commutation formulae
In the same vein, for any $n\geq 1$ we have
A proof for (123) and (124) is provided in Appendix E (on page E).
Using the commutation properties (123) and (124) for any $n\geq p$ we readily check that
with the backward semigroups
By (123) and (124) we also obtain
D.2 Some quantitative estimates
@@
Since $U$ is locally bounded with compact sub-level sets for any $\delta\geq 0$ and any sufficiently large $r>0$ we have
$e^{-\delta U}=e^{-\delta U}~1_{\{U>r\}}+e^{-\delta U}~1_{\{U\leq r\}}\leq\varsigma^{U}_{\delta}(r)\quad\mbox{\rm and}\quad\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}\geq 1/\varsigma^{U}_{\delta}(r)$
with the parameters
$\varsigma^{U}_{\delta}(r):=e^{-\delta r}+e^{\delta c_{U}(r)}\quad\mbox{\rm with}\quad c_{U}(r):=\sup_{x\in C_{U}(r)}|U(x)|\quad\mbox{\rm and}\quad C_{U}(r)=\{x~:~U(x)\leq r\}$
This yields
@@
We set $\sigma_{\beta}=\beta\sigma\beta^{\prime}$ and $\beta_{\tau}:=\tau^{-1}\beta$. With this notation at hand, after some elementary computations (cf. for instance Exemple 6.6 in [adm-24]) we have
@@ D.3 Lyapunov inequalities
${\cal S}_{2n}\left(\psi\right)/\varphi\leq c_{\delta}~e^{-{\cal U}_{\delta}}~\in{\cal B}_{0}(\mathbb{X})\quad\mbox{and}\quad{\cal S}_{2n-1}\left(\varphi\right)/\psi\leq c_{\delta}~e^{-{\cal V}_{\delta}}~\in{\cal B}_{0}(\mathbb{Y})$ (137)
@@
By (130) we have
${\cal S}_{1}(y,dx)$ $=$ $\frac{{\cal R}(y,dx)~e^{-U(x)}}{{\cal R}(e^{-U})(y)}\leq e^{W_{U}(y)}~{\cal R}(y,dx)~e^{-U(x)}$
This implies that
$e^{-\delta V(y)}{\cal S}_{1}\left(e^{\delta U}\right)(y)\leq e^{-(\delta V(y)-W_{U}(y))}~\|e^{-W}\|~\lambda_{(1-\delta)U}\left(1\right)$
Equivalently, some finite constant $c_{\delta}<\infty$ we have
${\cal S}_{1}\left(\varphi\right)/\psi\leq c_{\delta}~e^{-{\cal V}_{\delta}}$~
This ends the proof of the proposition.
D.4 Proof of Theorem 4.13
The Lyapunov inequalities (137) in Proposition D.9 ensure that the drift conditions stated in (60) hold with $n_{0}=1$ and $(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})$. Recall from the proof of Proposition D.5 that $e^{-\delta U}$ and $e^{-\delta V}$ are uniformly bounded for any $\delta\geq 0$. Since $\mathbb{X}$ and $\mathbb{Y}$ can be exhausted respectively by the compact sub-levels sets $C_{U}(r)$ and $C_{V}(r)$ of $U$ and $V$, there exists some $r_{1}>1$ such that $\lambda(C_{U}(r))\wedge\nu(C_{V}(r))>0$ for $r=r_{1}$ and thus for any $r\geq r_{1}$. This ensure there exists some $r_{0}>0$ such that for any $r\geq r_{0}$ we have
$\lambda(C_{\varphi}(r))\wedge\nu(C_{\psi}(r))>0.$
Following the proof of Proposition D.5, for $\lambda$-almost any $x\in C_{\varphi}(r)$ we have
@@
$\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\nu_{V}(e^{W_{U}})$
$\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\|e^{-\delta V}\|_{\mathbb{L}_{\infty}(\nu)}~\nu_{(1-\delta)V}\left(e^{W_{U}}\right)$
@@
$\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\lambda_{U}(e^{W^{V}})$
$\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}~\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)$
@@
On the other hand, using Lemma D.8 for any $x\in C_{\varphi}(r)$ and $r\geq r_{0}$ we arrive at
@@
Note that for any $r\geq r_{0}$ we have
@@
and this yields, for any $x\in C_{\varphi}(r)$ and $r\geq r_{0}$, the estimate
@@
[…]
@@
[…]
@@
where the last assertion comes from the fact that $\log{(1+u)}\leq u$, for any $1+u>0$. This implies that